OpenStudy (anonymous):

Solve for y. 8 and - 9 and = - 13 and=?

OpenStudy (anonymous):

So 8y-9y=-13 ? which would be -y = -13 or y = 13

OpenStudy (anonymous):

you can check by doing 8(13)-9(13) =

OpenStudy (anonymous):

8(13)-9(13) =(-13)